OpenStudy (nudeoua):

Solve for x: −3|2x + 6| = −12 (1 point) x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

OpenStudy (anonymous):

First you got to solve inside the brackets or the absolute value signs but since there are no like terms in there you now have to so you have to do distributive property and it will be -3(2x) + (-3)(6) = -6x + (-18) = -6x - 18 = -12. Now solve